# Heterogeneous Graph Neural Networks for Data-driven Traffic Assignment

### v1

@@ 3.3 Graph Edge Prediction
To predict the transportation performance at the edge level, the node embedding of the source node and destination node, and the real edge feature are concatenated and passed into the multi-layer perceptron to have the flow prediction $\tilde{f}_{e}$ of real link $e$:
$\tilde{f}_{e}=\texttt{MLP}([\bm{O}_{e,src}\oplus\bm{O}_{e,dst}\oplus\bm{X}_{e,r}];\bm{W}_{e},\bm{b}_{e}),$ (11)
where $\bm{O}_{e,src}$ and $\bm{O}_{e,dst}$ represent the source node embedding and destination node embedding, respectively. $\bm{W}_{e}$ and $\bm{b}_{e}$ is the learnable parameters associated with the multilayer perception. After generating the final prediction, selecting an appropriate loss function becomes crucial to ensure the model’s effective convergence. The proposed model utilizes a twofold loss function. The first part is the supervised loss, which measures the discrepancy between prediction and ground truth:
$L_{s}=\frac{1}{|\mathcal{E}_{r}|}\sum_{e\in\mathcal{E}_{r}}\|f_{e}-\tilde{f}_{e}\|,$ (12)
where the $f_{e}$ and $\tilde{f_{e}}$ represent the ground truth and prediction of flow on link $e\in\mathcal{E}$. The second part of the loss function is from the node-based flow conservation law, where the total flow of traffic entering a node equals the total flow of traffic exiting that node. The node-based flow conservation law can be represented mathematically:
@@
where $f_{ki}$ denotes the flow on the link $(k,i)$, $\Delta f_{i}$ represents the difference between flow receiving and sending at node $i$, $O_{v,i}$ represents the number of OD demand from $v$ to $i$. $\mathcal{V}$ and $\mathcal{V}_{OD}$ denote the entire node set and the origin-destination node set, respectively. The node-based flow conservation law can be considered as a normalization loss or auxiliary loss to ensure the predicted traffic flow at each node satisfies the flow conservation principle. One common way to incorporate conservation law into the loss function is to define a residual loss function:
$L_{f}=\sum_{i}\ |\sum_{k\in\mathcal{N}_{i}(i)}\tilde{f}_{ki}-\sum_{j\in\mathcal{N}_{o}(i)}\tilde{f}_{ij}-\Delta f_{i}|,$ (14)
where $\mathcal{N}_{i}(i)$ and $\mathcal{N}_{o}(i)$ represent the incoming edges and outcoming edges of node $i$, respectively. The normalization loss $L_{f}$ measures how the flow prediction satisfies the flow conservation law. Minimizing this loss function during training will encourage the model to learn traffic flow patterns that satisfy the conservation law. The total loss for the flow prediction $L_{total}$ is the weighted summation of the supervised loss and the conservation loss:
$L_{total}=w_{s}L_{s}+w_{f}L_{f},$ (15)
where the $w_{s}$ and $w_{f}$ represent the normalized weight for supervised loss and the conservation loss, respectively.
@@ 4 Numerical Experiments
To evaluate the accuracy, efficiency, and generalization capability of the proposed graph neural network, two numerical experiments are conducted. The first experiment is on urban transportation networks with synthetic traffic data. The second experiment is on multiple synthetic graphs with different topologies. Furthermore, we also synthesized variations in the OD demand and link capacity, as will be explained later.
4.1 Experiments on Urban Transportation Networks
As case studies, three urban transportation networks are selected: Sioux Falls network, East Massachusetts Network (EMA), and Anaheim network. The information about the network topology, link capacity, and the default OD demand of these networks are obtained from (Bar-Gera et al., 2023). The details of the network topologies are shown in Table 1. To create demand variation, we scaled the demand by a scaling factor according to
$\tilde{O}_{s,t}=\delta^{o}_{s,t}\ O_{s,t},$ (16)
$\tilde{c}_{a}=\delta^{c}_{a}\ c_{a},$ (17)
where $c_{a}$ is the original link capacity for link $a$, and $\delta^{c}_{a}$ is the scaling factor for link $a$. Capacity variations are considered to be due to traffic accidents, road construction/damage, and adverse weather conditions, which reduce the link capacity. In this work, three levels of capacity reduction are considered: (L): light disruption with $\delta^{c}_{a}\sim U(0.8,1.0)$; (M) moderate disruption with $\delta^{c}_{a}\sim U(0.5,1.0)$; (H) high disruption with $\delta^{c}_{a}\sim U(0.2,1.0)$. The size of the dataset for each network at each disruption scenario is 5000, which is split into the training set and the testing set with a ratio of 80% and 20%, respectively.
@@ 4.1 Experiments on Urban Transportation Networks
Figure 2: The illustrations of urban transportation networks, including Sioux Falls, EMA, and Anaheim. The link color represents the link capacity of each link.
The ground truth data used in training and testing steps is obtained from the user equilibrium (UE)-based traffic assignment. Specifically, the traffic flow and flow-capacity ratio (v/c ratio) of each link are the quantities of interest. The GNN model is implemented using PyTorch (Paszke et al., 2019), and DGL (Wang et al., 2019a). The preprocessing layer consists of a three-layer fully connected neural network with an embedding size of 64. The number of GNN layers in the proposed model is 3. The first layer is applied on virtual links, and the following layers are applied on real links. The number of heads in the attention block is 8. For hyper-parameter selection, the hidden layer size is chosen as 64, which is common in neural network implementation (Liu and Meidani, 2023b). Rectified Linear Unit (ReLU) activation (Agarap, 2018) is chosen to introduce nonlinearity. Mini-batch stochastic gradient descent was implemented in the training process. More specifically, the adaptive moment estimation (Adam) optimizer (Kingma and Ba, 2014) is adopted with a learning rate of 0.001. The batch size of training is 128. The weights for supervised loss and the conservation loss in equation 15 are chosen as 1.0 and 0.001, respectively, in order to keep both loss terms in the same order of magnitude. We evaluated the performance of our proposed heterogeneous GNN model (referred to by HetGAT) and compared it with three benchmark models: a fully connected neural network (FCNN), a homogeneous graph attention network (GAT), and a homogeneous graph convolution network (GCN). The FCNN consisted of five fully connected layers with an embedding size of 64. The GAT and GCN both have three layers of graph message passing layer, followed by three layers of FCNN with an embedding size of 64. The metrics to evaluate performance include the mean absolute error (MAE) and the normalized conversation loss $L_{f}$ given by
@@
$L_{f}=\frac{\sum_{i}|\sum_{k\in\mathcal{N}_{i}(i)}\tilde{f}_{ki}-\sum_{j\in\mathcal{N}_{o}(i)}\tilde{f}_{ij}-\Delta f_{i}|}{\sum_{s}\sum_{t}\tilde{O}_{s,t}},$ (19)
where $y$ and $\tilde{y}$ respectively represent the ground truth and predicted values for the flow-capacity ratio. Furthermore, we use the coefficient of determination, denoted by $R^{2}$ to measure the goodness-of-fit between the GNN prediction and the ground truth. The training histories of the studied models are shown in Figure 3. The results indicate that the FCNN model performed poorly during training compared to GNN-based models, as shown by the high training loss and early stagnation. In contrast, GCN and GAT models exhibited similar convergence rates. Our proposed model outperformed both GCN and GAT in terms of training loss. Especially when used for larger networks, the proposed model demonstrated superior convergence performance compared to GCN and GAT; for the Anaheim network, the training loss of the proposed model is almost 1/3 of that of GAT in the first 25 iterations. This is because GCN and GAT only consider homogeneous edges, which limits the message passing to adjacent nodes. In contrast, the proposed GNN model uses virtual links and provides augmented connectivity to long-hop node pairs, which makes the node feature updating in HetGAT more efficiently.
@@
After the training is finished, the model performance is evaluated on the testing set. The experiments under the urban road network are conducted in three different settings. The first setting, referred to as LMH-LMH, involved using all levels of disruption (flow reduction scaling levels) in both the training and testing sets. The second setting, namely L-M, involves training the model on light disruption data and testing it on medium disruption data. The third case, which is labeled as M-H, involves training on medium disruption data and testing on high disruption data. The L-M and M-H scenarios will therefore involve unseen cases that don’t exist in the training data. Figure 4 plotted the predicted value and ground truth of the flow-capacity ratio on multiple samples in the Anaheim network under the LMH-LMH setting. In total, the flow-capacity ratio on 10,000 edges is predicted using HetGAT, GAT, and GCN, respectively. Figure 4 indicates that HetGAT has a relatively higher correlation coefficient, which outperforms GAT and GCN. Table 2 summarizes the prediction performance of all methods under different settings, and shows that HetGAT, compared to other models, offers better performance. When the graph size increases, the proposed model can maintain a relatively low MAE compared to GCN and GAT. For instance, in the Anaheim network, HetGAT offers MAE values that are 39%, 19%, and 31% lower than GAT values, in LMH-LMH, L-M, and M-H settings, respectively. Also, the conservation errors of HetGAT are 52%, 29%, and 55% smaller in LMH-LMH, L-M, and M-H settings, respectively. This shows that the inclusion of virtual links can assist GNN models in better learning the traffic flow patterns.
@@
Table 2: Performance comparison of HetGAT with benchmark methods. Three different settings are included: LMH-LMH, L-M, M-H. The mean absolute error, normalized conservation loss, and correlation coefficient are used to evaluate the prediction performance on the testing set.
@@
MAE $L_{f}$ $R^{2}$ MAE $L_{f}$ $R^{2}$ MAE $L_{f}$ $R^{2}$
Sioux Falls FCNN 0.158 0.324 0.728 0.139 0.062 0.812 0.215 0.238 0.643
GAT 0.046 0.092 0.956 0.022 0.018 0.991 0.085 0.057 0.889
GCN 0.046 0.096 0.955 0.023 0.019 0.990 0.086 0.060 0.889
HetGAT 0.028 0.054 0.985 0.026 0.018 0.986 0.062 0.039 0.937
EMA FCNN 0.238 0.422 0.621 0.225 0.144 0.795 0.263 0.446 0.771
GAT 0.063 0.131 0.952 0.046 0.033 0.991 0.132 0.090 0.842
GCN 0.071 0.167 0.944 0.034 0.043 0.986 0.161 0.135 0.814
HetGAT 0.038 0.090 0.982 0.034 0.027 0.989 0.090 0.055 0.921
Anaheim FCNN 0.111 0.721 0.822 0.199 0.237 0.801 0.186 0.386 0.840
GAT 0.051 0.263 0.951 0.042 0.081 0.958 0.090 0.184 0.881
GCN 0.072 0.402 0.903 0.062 0.109 0.911 0.116 0.245 0.827
HetGAT 0.031 0.126 0.981 0.034 0.059 0.979 0.062 0.082 0.939
@@
As an additional experiment, we consider a realistic scenario where the demand values are not known for the entire OD pairs. To evaluate the accuracy and robustness of the proposed model under incomplete OD demand scenarios, we introduce a random mask to the original OD demand, to simulate these missing demand scenarios. Specifically, given a specific missing ratio, we randomly select a number of OD pairs and mask their corresponding OD demand values as zeros in the input node feature. This way, the model is expected to learn the inherent patterns and structures of the traffic network, even when some of the demand information is missing. The effectiveness of the proposed model under incomplete OD demand scenarios will be evaluated by comparing the predicted traffic flows against the ground truth data obtained from the complete OD demand. Three missing ratios are considered in the experiment: 20%, 30%, and 40%. The training setting and the hyperparameters remain the same as those in the aforementioned experiments. Table 3 summarizes the results of prediction performance under different missing rate scenarios for the LMH-LMH setting. The FCNN model is not considered in these experiments because of their very poor performance in the previous experiments under full OD demand. As expected, overall the performances under incomplete OD demand drop, compared to that in the complete demand scenario, reported in Table 2. However, HetGAT still outperforms GAT and GCN under different networks and different missing ratios. In particular, the MAE values are about 50% smaller to those offered by GAT and GCN under various missing ratios. Also, the flow predictions by HetGAT have better compliance with the flow conservation law compared with GAT and GCN.
@@
Table 3: Comparison of the performance of HetGAT with that of GAT and GCN under incomplete OD demand. Three missing ratios are considered in the experiments: 20%, 30%, and 40%. The mean absolute error (MAE), normalized conservation loss ($L_{f}$), and the coefficient of determination, $R^{2}$, are used to evaluate the prediction performance on the testing set.
Network Model Missing Ratio
$20\%$ $30\%$ $40\%$
MAE $L_{f}$ $R^{2}$ MAE $L_{f}$ $R^{2}$ MAE $L_{f}$ $R^{2}$
Sioux Falls GAT 0.048 0.098 0.954 0.049 0.102 0.953 0.049 0.099 0.952
GCN 0.049 0.102 0.951 0.049 0.101 0.953 0.049 0.101 0.952
HetGAT 0.033 0.068 0.981 0.034 0.069 0.979 0.033 0.068 0.982
EMA GAT 0.068 0.145 0.938 0.070 0.149 0.931 0.070 0.149 0.933
GCN 0.079 0.200 0.936 0.079 0.204 0.935 0.077 0.188 0.939
HetGAT 0.046 0.111 0.970 0.048 0.115 0.970 0.048 0.105 0.972
Anaheim GAT 0.067 0.152 0.924 0.061 0.135 0.933 0.071 0.156 0.914
GCN 0.105 0.225 0.835 0.109 0.237 0.802 0.106 0.239 0.827
HetGAT 0.035 0.060 0.975 0.034 0.069 0.969 0.034 0.058 0.977

### v2

@@ 3.3 Graph Edge Prediction
To predict the traffic flow at the edge level, the node embedding of the source node and destination node, and the normalized real edge feature are concatenated and passed through a feed-forward neural network. In this paper, we consider the flow-capacity ratio $\tilde{\alpha}_{e}$ as the quantity of the final link prediction, which is the link flow normalized by the link capacity:
$\tilde{\alpha}_{e}=\texttt{MLP}([\bm{o}_{u}\oplus\bm{o}_{v}\oplus\bm{\beta}^{r}_{e}];\bm{W}_{o},\bm{b}_{o}),\quad\forall e=(u,v)\in\mathcal{E}_{r}$ (10)
where $\bm{o}_{(\cdot)}$ represent the node embedding. $\bm{W}_{o}$ and $\bm{b}_{o}$ is the learnable parameters associated with the multilayer perception. The predicted link flows $\tilde{f}_{e}$ can be calculated by multiplying the link capacity with the predicted flow-capacity ratio. Subsequently, selecting an appropriate loss function becomes crucial to ensure the model’s effective convergence. The proposed model employs a composite loss function comprising two components. The first part is the supervised loss, which measures the difference between prediction and ground truth. It considers both the discrepancy from the flow-capacity ratio $L_{\alpha}$ and the link flow $L_{f}$ on each link:
$\begin{split}L_{s}&=L_{\alpha}+L_{f}\\ &=\frac{1}{|\mathcal{E}_{r}|}\sum_{e\in\mathcal{E}_{r}}\|{\alpha}_{e}-\tilde{\alpha}_{e}\|+\frac{1}{|\mathcal{E}_{r}|}\sum_{e\in\mathcal{E}_{r}}\|f_{e}-\tilde{f}_{e}\|,\end{split}$ (11)
where the ${\alpha}_{e}$ and $\tilde{\alpha}_{e}$ represent the ground truth and prediction of flow-capacity ratio on link $e\in\mathcal{E}$. The $f_{e}$ and $\tilde{f_{e}}$ represent the ground truth and prediction of link flow on link $e\in\mathcal{E}$. The second part of the loss function originates from the principle of node-based flow conservation, where the total flow of traffic entering a node equals the total flow of traffic exiting that node. The node-based flow conservation law can be represented mathematically:
@@
where $f_{ki}$ denotes the flow on the link $(k,i)$, $\Delta f_{i}$ represents the difference between flow receiving and sending at node $i$, $O_{v,i}$ represents the number of OD demand from $v$ to $i$. $\mathcal{V}_{OD}$ denote the origin-destination node set. The node-based flow conservation law can be considered as a normalization loss, thereby ensuring compliance with the fundamental principle of flow conservation. One common way to incorporate conservation law into the loss function is to define a residual loss function:
$L_{c}=\sum_{i}\ |\sum_{k\in\mathcal{N}_{i}(i)}\tilde{f}_{ki}-\sum_{j\in\mathcal{N}_{o}(i)}\tilde{f}_{ij}-\Delta f_{i}|,$ (13)
where $\mathcal{N}_{i}(i)$ represent the incoming edges of node $i$. The normalization loss $L_{c}$ measures how the flow prediction satisfies the flow conservation law. Minimizing this loss function during training will encourage the model to learn traffic flow patterns that satisfy the conservation law. Consequently. the total loss for the flow prediction $L_{total}$ is the weighted summation of the supervised loss and the conservation loss:
$L_{total}=w_{\alpha}L_{\alpha}+w_{f}L_{f}+w_{c}L_{c},$ (14)
where the $w_{\alpha}$, $w_{f}$ and $w_{c}$ represent the normalized weight for supervised loss of flow-capacity ratio, supervised loss of actual flow, and the conservation loss, respectively.
@@ 4 Numerical Experiments
Two numerical experiments are conducted to evaluate the accuracy, efficiency, and generalization capability of the proposed graph neural network. The first experiment is on urban transportation networks. The second experiment is on multiple synthetic graphs with different topologies. The details of the experiments will be explained in the following sections.
@@
4.1 Experiments on Urban Transportation Networks
As case studies, three urban transportation networks are selected: Sioux Falls network, East Massachusetts Network (EMA), and Anaheim network. The information about the network topology, link characteristics, and the OD demand of these networks are obtained from (Bar-Gera et al., 2023). The statistics and the illustration of the network topologies are shown in Table 1 and Figure 2, respectively. To create demand variation, we scaled the demand by a scaling factor according to
$\tilde{O}_{s,t}=\delta^{o}_{s,t}\ O_{s,t},$ (15)
$\tilde{c}_{a}=\delta^{c}_{a}\ c_{a},$ (16)
where $c_{a}$ is the original link capacity for link $a$, and $\delta^{c}_{a}$ is the scaling factor for link $a$. Capacity variations are considered to be due to traffic accidents, road construction/damage, and adverse weather conditions, which reduce the link capacity. In this work, three levels of capacity reduction are considered: (L): light disruption with $\delta^{c}_{a}\sim U(0.8,1.0)$; (M) moderate disruption with $\delta^{c}_{a}\sim U(0.5,1.0)$; (H) high disruption with $\delta^{c}_{a}\sim U(0.2,1.0)$.
The size of the dataset for each network at each disruption scenario is 5000, which is split into the training set and the testing set with a ratio of 80% and 20%, respectively. To demonstrate the dataset is sufficiently diverse to cover enough scenarios, the coefficient of variation of network link capacity and OD demand is calculated and the histogram of the link capacity and the OD demand of training and testing data is shown in Figure 3. The OD demand of each network is normalized to 100 in order to facilitate a standardized comparison across different networks regardless of their actual size or demand volumes. The minimum coefficient of variation of link capacity and OD demand among the three networks is 0.45 and 0.22, respectively, which indicates the training and testing data are sufficiently diverse to cover different scenarios (Bedeian and Mossholder, 2000; Campbell et al., 2010).
@@ 4.1 Experiments on Urban Transportation Networks
Figure 2: The illustrations of urban transportation networks, including Sioux Falls, EMA, and Anaheim.
Figure 3: The histogram of the link capacity and OD demand in the training and testing data. Three transportation networks are considered including Sioux Falls, EMA, and Anaheim network.
The training and testing dataset are obtained by solving UE-TAP with the Frank-Wolfe algorithm (Fukushima, 1984). The algorithm converges when the square root of the sum of the squared differences between the link flow in two successive iterations, normalized by the sum of the values of the link flow, falls below the threshold of 1e-5. The GNN model is implemented using PyTorch (Paszke et al., 2019) and DGL (Wang et al., 2019a). The preprocessing layer consists of a three-layer fully connected neural network with an embedding size of 32. The number of GNN layers in the proposed model is 4, including two V-Encoders and two R-Encoders. The number of heads in the attention block is 8. For hyper-parameter selection, the hidden layer size is chosen as 64, which is common in neural network implementation (Liu and Meidani, 2023c, b). The learning rate and batch size of training are 0.001 and 128, respectively. The weights of $L_{\alpha}$, $L_{f}$, and $L_{c}$ in equation 14 are chosen as 1.0, 0.005, and 0.05, respectively.
We evaluated the performance of our proposed heterogeneous GNN model (referred to by HetGAT) and compared it with three benchmark models: a fully connected neural network (FCNN), a homogeneous graph attention network (GAT), and a homogeneous graph convolution network (GCN). The FCNN consisted of five fully connected layers with an embedding size of 64. The GAT and GCN both have four layers of graph message passing layer, followed by three layers of FCNN with an embedding size of 64. The metrics to evaluate performance include the mean absolute error (MAE), root mean square error (RMSE), and the normalized conversation loss $\tilde{L}_{c}$:
@@
$\mathrm{RMSE}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}(y_{i}-\tilde{y}_{i})^{2}},$ (18)
$\tilde{L}_{c}=\frac{\sum_{i}|\sum_{k\in\mathcal{N}_{i}(i)}\tilde{f}_{ki}-\sum_{j\in\mathcal{N}_{o}(i)}\tilde{f}_{ij}-\Delta f_{i}|}{\sum_{s}\sum_{t}\tilde{O}_{s,t}},$ (19)
where $y$ and $\tilde{y}$ respectively represent the ground truth and predicted values for quantity of interest. We conducted a 5-fold cross-validation for each experiment to ensure the robustness of our results across different subsets of the data.
The training histories of the studied models are shown in Figure 4. The results indicate that the FCNN model performed poorly during training compared to GNN-based models, as shown by the high training loss and early stagnation. In contrast, GCN and GAT models exhibited similar convergence rates. Our proposed model outperformed both GCN and GAT in terms of training loss. Especially when used for larger networks, the proposed model demonstrated superior convergence performance compared to GCN and GAT; for the Anaheim network, the training loss of the proposed model is almost 1/3 of that of GAT in the first 25 iterations. This is because GCN and GAT only consider homogeneous edges, which limits the message passing to adjacent nodes. In contrast, the proposed GNN model uses virtual links and provides augmented connectivity to long-hop node pairs, which makes the node feature updating in HetGAT more efficient.
@@
(a) Sioux Falls network
(b) EMA network
(c) Anaheim network
Figure 5: Illustrations of the link-wise flow distribution for different transportation networks. Three transportation networks under major disruption are considered, including Sioux Falls, EMA, and Anaheim networks. 50 links are selected for each network.
After the training is finished, the model performance is evaluated on the testing set. The experiments under the urban road network are conducted in three different settings. The first setting, referred to as LMH-LMH, involved using all levels of disruption (flow reduction scaling levels) in both the training and testing sets. The second setting, namely L-M, involves training the model on light disruption data and testing it on medium disruption data. The third case, which is labeled as M-H, involves training on medium disruption data and testing on high disruption data. The L-M and M-H scenarios will therefore involve unseen cases that don’t exist in the training data. Figure 5 visualizes the variations in link-wise flow in Sioux Falls, EMA, and Anaheim networks under high disruption. The average coefficients of variation of actual link flow for three networks are 0.297, 0.242, and 0.201, respectively. Figure 6 plotted the predicted value and ground truth of the flow-capacity ratio on multiple samples in the Anaheim network under the LMH-LMH setting. In total, the flow-capacity ratio on 10,000 edges is predicted using HetGAT, GAT, and GCN, respectively. Figure 6 indicates that HetGAT has a relatively higher correlation coefficient, which outperforms GAT and GCN.
Table 2 summarizes the prediction performance of all methods under different settings, and shows that HetGAT, compared to other models, offers better performance. When the graph size increases, the proposed model maintains a relatively low MAE compared to GCN and GAT. For instance, in the EMA network, HetGAT offers flow MAEs that are 39.5%, 56.4%, and 38.8% lower than the second best result, in LMH-LMH, L-M, and M-H settings, respectively. In the Anaheim network, HetGAT offers flow MAEs that are 27.2%, 47.4%, and 33.1% lower than the second-best result in LMH-LMH, L-M, and M-H settings, respectively. This shows that the inclusion of virtual links can assist GNN models in better learning the traffic flow patterns. In addition to the prediction accuracy, the training time is also an important factor in evaluating the efficiency and practicality of machine learning models. The computational time is mainly threefold: time for solving UE-TAP, GNN training time, and GNN inference time. For Sioux Falls, EMA, and Anaheim networks, the time for solving UE-TAP is 56.8, 575.2, and 2612.5 min, respectively. In comparison, the training time of HetGAT is 26.8, 28.9, and 59.7 min, respectively. Moreover, for every 1000 graphs, the inference time of the proposed model is notably efficient at 0.13, 0.15, and 0.31 min, respectively. The computational time of each component demonstrates the computational efficiency of the proposed model for solving traffic assignment problems.
@@
Table 2: Performance comparison of HetGAT with benchmark methods. Three different settings are included: LMH-LMH, L-M, M-H. The mean absolute error, root mean square error, and normalized conservation loss are used to evaluate the prediction performance on the testing set.
@@
Flow
($1\times 10^{2}$)
Link
utilization
$\tilde{L}_{c}$
Flow
($1\times 10^{2}$)
Link
utilization
$\tilde{L}_{c}$
Flow
($1\times 10^{2}$)
Link
utilization
$\tilde{L}_{c}$
MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE
Sioux Falls FCNN 19.31 30.22 16.29% 26.39% 0.38 18.64 27.31 15.87% 24.55% 0.37 32.85 48.08 30.09% 44.99% 0.70
GCN 5.62 8.80 4.78% 7.74% 0.11 5.43 7.95 4.62% 7.15% 0.14 9.57 14.00 8.64% 12.87% 0.22
GAT 5.60 8.75 4.74% 7.68% 0.11 5.54 8.18 4.62% 7.15% 0.11 9.55 13.99 8.76% 13.10% 0.20
HetGAT 3.58 5.56 2.99% 4.74% 0.07 3.73 5.39 2.99% 4.36% 0.07 7.01 10.39 6.32% 9.52% 0.08
EMA FCNN 7.51 14.82 25.56% 63.89% 0.81 9.04 17.70 32.36% 61.87% 0.46 13.82 24.68 60.47% 122.94% 0.62
GCN 1.99 3.92 7.62% 18.32% 0.17 2.39 4.69 7.78% 14.65% 0.08 3.66 6.53 14.35% 30.33% 0.12
GAT 1.77 3.62 6.77% 16.91% 0.21 2.39 4.69 8.57% 16.38% 0.12 3.83 6.87 16.01% 32.54% 0.17
HetGAT 0.98 2.21 3.48% 8.08% 0.07 1.05 2.06 3.52% 6.85% 0.03 2.49 4.79 9.24% 20.50% 0.07
Anaheim FCNN 10.61 16.73 24.72% 45.03% 0.81 15.82 28.59 38.44% 69.66% 0.23 18.74 29.82 49.70% 84.45% 0.25
GCN 2.18 3.44 6.54% 11.53% 0.29 3.25 5.88 9.31% 17.09% 0.07 3.85 6.13 12.37% 20.74% 0.08
GAT 1.47 2.55 4.77% 8.69% 0.16 2.32 4.18 7.42% 13.45% 0.05 3.31 5.14 9.59% 16.30% 0.05
HetGAT 1.04 1.81 2.97% 5.46% 0.08 1.19 1.85 3.31% 5.45% 0.02 2.17 3.37 6.66% 10.85% 0.03
@@
As an additional experiment, we consider a realistic scenario where the regional OD demand values are incomplete by introducing a random binary mask to the original OD demand, More specifically, given a specific missing ratio, we randomly select a number of OD pairs and mask their corresponding OD demand values as zeros in the input node feature. In this way, the model is expected to learn the inherent patterns and structures of the transportation network, even when some of the demand information is missing. The effectiveness of the proposed model under incomplete OD demand scenarios will be evaluated by comparing the predicted traffic flows against the ground truth data obtained from the complete OD demand. Three missing ratios are considered in the experiment: 20%, 30%, and 40%. The training setting and the hyperparameters remain the same as those in the aforementioned experiments. Table 3 summarizes the results of prediction performance under different missing rate scenarios for the LMH-LMH setting. The FCNN model is not considered in these experiments because of their very poor performance in the previous experiments under full OD demand. HetGAT still outperforms GAT and GCN under different networks and different missing ratios. additionally, the flow predictions by HetGAT have relatively better compliance with the flow conservation law compared with GAT and GCN.
@@
Table 3: Comparison of the performance of HetGAT with benchmark under incomplete OD demand. Three missing ratios are considered in the experiments: 20%, 30%, and 40%. The mean absolute error, root mean square error, and normalized conservation loss, are used to evaluate the prediction performance on the testing set.
Network Model Missing ratio = 20% Missing ratio = 30% Missing ratio =40%
Flow
($1\times 10^{2}$)
Link
utilization
$\tilde{L}_{c}$
Flow
($1\times 10^{2}$)
Link
utilization
$\tilde{L}_{c}$
Flow
($1\times 10^{2}$)
Link
utilization
$\tilde{L}_{c}$
MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE
Sioux Falls GCN 6.50 10.00 5.39% 8.62% 0.15 6.54 10.02 5.72% 9.19% 0.15 7.54 11.06 6.68% 10.60% 0.16
GAT 6.64 10.11 5.83% 9.25% 0.14 7.19 10.50 6.11% 9.45% 0.14 6.68 10.13 5.78% 9.20% 0.15
HetGAT 4.05 6.04 3.33% 5.06% 0.08 4.13 6.14 3.38% 5.13% 0.09 4.15 6.08 3.45% 5.16% 0.09
EMA GCN 2.14 4.11 8.06% 16.94% 0.24 2.15 4.10 7.87% 16.47% 0.25 2.14 4.12 8.03% 17.09% 0.26
GAT 1.87 3.71 6.89% 16.65% 0.18 1.86 3.73 7.00% 16.77% 0.19 1.87 3.73 7.07% 16.97% 0.19
HetGAT 1.15 2.37 3.98% 8.40% 0.08 1.15 2.37 4.00% 8.24% 0.08 1.21 2.41 4.17% 8.67% 0.09
Anaheim GCN 2.16 3.38 6.31% 10.86% 0.14 2.24 3.43 6.65% 11.18% 0.14 2.28 3.56 6.82% 11.72% 0.15
GAT 1.62 2.70 4.80% 8.70% 0.12 1.74 2.85 5.21% 9.34% 0.13 1.68 2.78 4.98% 8.97% 0.12
HetGAT 1.11 1.91 3.17% 5.78% 0.06 1.08 1.87 3.08% 5.65% 0.05 1.09 1.89 3.12% 5.72% 0.05